Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $80 per unit. Variable expenses are $40.00 per unit, and fixed expenses total $180,000 per year. Its operating results for last year were as follows: Sales $ 2,080,000 Variable expenses 1,040,000 Contribution margin 1,040,000 Fixed expenses 180,000 Net operating income $ 860,000 Required: Answer each question independently based on the original data:

 1. What is the product's CM ratio?

 2. Use the CM ratio to determine the break-even point in dollar sales.

 3. If this year's sales increase by $52,000 and fixed expenses do not change, how much will net operating income increase?

 4-a. What is the degree of operating leverage based on last year's sales?

 4-b. Assume the president expects this year's sales to increase by 11%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

 5. The sales manager is convinced that a 15% reduction in the selling price, combined with a $77,000 increase in advertising, would increase this year's unit sales by 25%. 

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented? 

b. Do you recommend implementing the sales manager's suggestions?

 6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.40 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $860,000 net operating income as last year?
Answer #1

Answer

  • All working forms part of the answer and explained through ‘working’ equation column.
  • General working for last year’s data

Amount

per unit

Units

A

Sales

2080000

80

26000

B

variable expenses

1040000

40

26000

C=A-B

Contribution margin

1040000

40

26000

D

Fixed expenses

180000

E=C-D

Net Operating Income

860000

  • Requirement 1: Products CM ratio

A

Contribution Margin

1040000

B

Sales

2080000

C=A/B

CM ratio

50%

  • Requirement 2: Break Even point in Sales dollars

A

Fixed Cost

180000

B

CM ratio

50%

C=A/B

Break Even point in dollar Sales

360000

  • Requirement 3: Change in Net operating income if sales increased by $52000

A

Sales increase by

52000

B

CM ratio

50%

C=A x B

Net operating income will increase y

26000

  • Requirement 4(a): Degree of Operating leverage

A

Contribution margin

1040000

B

Net Operating Income

860000

C=A/B

Operating leverage (in times)

1.209302326

  • Requirement 4(b): % change in Net operating Income

A

Sales increase by

228800

B

CM Ratio

50%

C=A x B

Contribution margin

114400

D

Operating leverage

1.209302326

E=C/D

Increase in Net operating income

94600

F

Old Net Operating Income

860000

G=E/F

% Change

11%

  • Requirement 5: What if Unit sales increased by 25% along with advertising cost

Amount

per unit

Units

A

Sales

2210000

68

32500

B

variable expenses

1300000

40

32500

C=A-B

Contribution margin

910000

28

32500

D

Fixed expenses

180000

E

Advertising

77000

F=C-D-E

Net Operating Income

653000

---Since Net operating income has decreased from $860000 to $653000, Sales manager’s suggestion should not be accepted.

  • Requirement 6: President’s View

Amount

per unit

Units

A

Sales

2600000

80

32500

B

variable expenses

1300000

40

32500

C

Commission

78000

2.4

32500

D=A-B-C

Contribution margin

1222000

37.6

32500

E

Fixed expenses

180000

F

Advertising

x

G=D-E-F

Net Operating Income

860000

Hence, the equation will be ---: D – E – F = G
1222000 – 180000 – x = 860000
1222000 – 180000 – 860000 = x
x = 182000

Therefore, the advertising expense can be increased to $182000 to keep the Net operating Income same as last year’s at $860000